# M3-Embedding: Multi-Linguality, Multi-Functionality, Multi-Granularity Text Embeddings Through Self-Knowledge Distillation

### v4

@@ 4.2 Cross-Lingual Retrieval
We make evaluation for the cross-lingual retrieval performance with the MKQA benchmark Longpre et al. (2021), which includes queries in 25 non-English languages. For each query, it needs to retrieve the passages containing answers from the English Wikipedia corpus. In our experiment, we make use of the well-processed corpus offered by the BEIR [^6] https://huggingface.co/datasets/BeIR/nq Thakur et al. (2021). Following the previous study Izacard et al. (2022), we report Recall@100 as the primary metric (Recall@20 is reported as an auxiliary metric in the Appendix C.1). For Dense+Sparse method and All method, we set the same weights as in MIRACL dataset.
@@ 4.3 Multilingual Long-Doc Retrieval
We evaluate the retrieval performance with longer sequences with two benchmarks: MLDR (Multilingual Long-Doc Retrieval), which is curated by the multilingual articles from Wikipedia, Wudao and mC4 (see Table 7), and NarrativeQA Kočiský et al. (2018); Günther et al. (2023), which is only for English. In addition to the previous baselines, we further introduce JinaEmbeddingv2 Günther et al. (2023), text-embedding-ada-002 and text-embedding-3-large from OpenAI given their outstanding long-doc retrieval capability. For Dense+Sparse method, we set $w_{1}=0.2$, $w_{2}=0.8$ and $w_{3}=0$ in equation(1). For All method, we set $w_{1}=0.15$, $w_{2}=0.5$ and $w_{3}=0.35$ in equation(1).
@@ 4.4 Ablation study
Impact of multi-stage training. We also make explorations for the impacts from different training stages. Fine-tuning indicates the direct fine-tuning from XLM-RoBERTA Conneau et al. (2020); RetroMAE+Fine-tuning refers to the fine-tuning on the pre-trained model from RetroMAE Xiao et al. (2022). Meanwhile, RetroMAE+Unsup+Fine-tuning involves fine-tuning on a model that is trained with RetroMAE and then pre-trained on unsupervised data. The results are presented in Table 6. We can observe that RetroMAE can significantly improve the retrieval performance, and pre-training on unsupervised data can further enhance the retrieval quality of the embedding model. (More detailed results are available in Appendix C.1.)
@@ B.1 Experimental Hyperparameters
We adopt a further pre-trained XLM-RoBERTa [^8] https://huggingface.co/FacebookAI/xlm-roberta-large as the foundational model. We extend the max position to 8192 and update the model via the RetroMAE Xiao et al. (2022) method. The data comprises Pile Gao et al. (2020), Wudao Yuan et al. (2021), and mC4 Raffel et al. (2020) datasets. We sampled a total of 184 million text samples from these sources, covering 105 languages. The maximum sequence length is 8192 and the learning rate is $7\times 10^{-5}$. The batch size is set to 32 and we accumulate the gradient over 16 steps. Pre-training is conducted on 32 A100(40GB) GPUs for 20,000 steps.

### v5

@@ 4.2 Cross-Lingual Retrieval
We make evaluation for the cross-lingual retrieval performance with the MKQA benchmark longpre-etal-2021-mkqa, which includes queries in 25 non-English languages. For each query, it needs to retrieve the passages containing answers from the English Wikipedia corpus. In our experiment, we make use of the well-processed corpus offered by the BEIR [^6] https://huggingface.co/datasets/BeIR/nq thakur2021beir. Following the previous study mcontriever, we report Recall@100 as the primary metric (Recall@20 is reported as an auxiliary metric in the Appendix C.1). For Dense+Sparse method and All method, we set the same weights as in MIRACL dataset.
@@ 4.3 Multilingual Long-Doc Retrieval
We evaluate the retrieval performance with longer sequences with two benchmarks: MLDR (Multilingual Long-Doc Retrieval), which is curated by the multilingual articles from Wikipedia, Wudao and mC4 (see Table 7), and NarrativeQA kocisky-etal-2018-narrativeqa; jina, which is only for English. In addition to the previous baselines, we further introduce JinaEmbeddingv2 jina, text-embedding-ada-002 and text-embedding-3-large from OpenAI given their outstanding long-doc retrieval capability. For Dense+Sparse method, we set $w_{1}=0.2$, $w_{2}=0.8$ and $w_{3}=0$ in equation(1). For All method, we set $w_{1}=0.15$, $w_{2}=0.5$ and $w_{3}=0.35$ in equation(1).
@@ 4.4 Ablation study
Impact of multi-stage training. We also make explorations for the impacts from different training stages. Fine-tuning indicates the direct fine-tuning from XLM-RoBERTA xlm-roberta; RetroMAE+Fine-tuning refers to the fine-tuning on the pre-trained model from RetroMAE xiao-etal-2022-retromae. Meanwhile, RetroMAE+Unsup+Fine-tuning involves fine-tuning on a model that is trained with RetroMAE and then pre-trained on unsupervised data. The results are presented in Table 6. We can observe that RetroMAE can significantly improve the retrieval performance, and pre-training on unsupervised data can further enhance the retrieval quality of the embedding model. (More detailed results are available in Appendix C.1.)
@@ Acknowledgements
Language Source #train #dev #test #cropus Avg. Length of Docs
ar Wikipedia 1,817 200 200 7,607 9,428
de Wikipedia, mC4 1,847 200 200 10,000 9,039
en Wikipedia 10,000 200 800 200,000 3,308
es Wikipedia, mC4 2,254 200 200 9,551 8,771
fr Wikipedia 1,608 200 200 10,000 9,659
hi Wikipedia 1,618 200 200 3,806 5,555
it Wikipedia 2,151 200 200 10,000 9,195
ja Wikipedia 2,262 200 200 10,000 9,297
ko Wikipedia 2,198 200 200 6,176 7,832
pt Wikipedia 1,845 200 200 6,569 7,922
ru Wikipedia 1,864 200 200 10,000 9,723
th mC4 1,970 200 200 10,000 8,089
zh Wikipedia, Wudao 10,000 200 800 200,000 4,249
Total 41,434 2,600 3,800 493,709 4,737
Table 7: Specifications of MultiLongDoc dataset.
@@ B.1 Experimental Hyperparameters
We adopt a further pre-trained XLM-RoBERTa [^8] https://huggingface.co/FacebookAI/xlm-roberta-large as the foundational model. We extend the max position to 8192 and update the model via the RetroMAE xiao-etal-2022-retromae method. The data comprises Pile pile, Wudao yuan2021wudaocorpora, and mC4 2019t5 datasets. We sampled a total of 184 million text samples from these sources, covering 105 languages. The maximum sequence length is 8192 and the learning rate is $7\times 10^{-5}$. The batch size is set to 32 and we accumulate the gradient over 16 steps. Pre-training is conducted on 32 A100(40GB) GPUs for 20,000 steps.